OpenStudy (anonymous):

what is (a+b)^9 / (a+b)^4

OpenStudy (shamim):

x^9/x^4=?

OpenStudy (anonymous):

(a+b)^5 because when you divide expressions, and they have the same base or variables, you subtract the exponents

OpenStudy (shamim):

x^(9-4)=x^5 right?

OpenStudy (anonymous):

yeah that makes sense